crystal sells coffee. her special blend is a combination of french roast @ $7.50 a lb. and kona deluxe @ $11.50 a lb. how many pounds of kona should she mix with 12 lbs. of french roast to get a blend to be sold for $10.00 a lb.?
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

7.5*12+11.5X=10(12+X)
90+11.5X=120+10X
11.5X-10X=120-90
1.5X=30
X=30/1.5
X=20 LBS OF KONA IS NEEDED.
PROOF
7.5*12+11.5*20=10(12+20)
90+230=10(32)
320=320

Answer by stanbon(57289)

crystal sells coffee. her special blend is a combination of french roast @ $7.50 a lb. and kona deluxe @ $11.50 a lb. how many pounds of kona should she mix with 12 lbs. of french roast to get a blend to be sold for $10.00 a lb.?
--------------
Amt. of french is 12 lbs; value of this is 7.50*12 = 90 dollars
------------
Let amt of kona be x lbs; value of this is 11.50x dollars
---------------
Amt of mixture is 12+x lbs; value is 10(12+x)=120+10x dollars
----------------------
EQUATION:
90 + 11.50x = 120+10x
1.5x = 30
x = 20 lbs. (Amount of kona that must be added to make the blend)
============
Cheers,
Stan H.